OpenStudy (anonymous):

The results of a fitness trial is a random variable X which is normally distributed with mean µ and standard deviation 2.4. A researcher uses the results from a random sample of 90 trials to calculate a 98% confidence interval for µ. What is the width of this interval ?

OpenStudy (anonymous):

@phi pls help

OpenStudy (phi):

if you have a random sample of 90 trials and find the average of them, you will get mu \( \pm\) some standard deviation

OpenStudy (anonymous):

how to find the average ?

OpenStudy (phi):

the std dev would be the population std dev / sqrt(size of sample) \[ \sigma= \frac{\sigma_{pop}} { \sqrt{90} } =\frac{2.4} { \sqrt{90} } \]

OpenStudy (phi):

you don't find the average... you are told it is \( \mu\) what they want are the lower and upper bounds x so that \[ \mu \pm x\] represents 90% of the area under the curve.

OpenStudy (phi):

**98%

OpenStudy (anonymous):

we dont have to find mu ?

OpenStudy (phi):

no. find the number of std deviations so that you get 98% of the area